The average number of milligrams (mg) of cholesterol in a cup of certain brand of ice cream is 660 mg, and the standard deviation is 35 mg. assume the variable is normally distrubuted.


If a sample of 10 cups of ice cream is selected, what is the probability that the mean of the sample will be larger than 670 mg


1
Expert's answer
2021-04-29T14:45:39-0400

Let "X="  the mean of the sample: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=660\ mg, \sigma=35\ mg, n=10"


"P(X>670)=1-P(X\leq670)"

"=1-P(Z\leq\dfrac{670-660}{35/\sqrt{10}})"

"\approx1-P(Z\leq0.9035)\approx0.1831"

The probability that the mean of the sample will be larger than 670 mg is 0.1831.